# Interactive Data Harmonization with LLM Agents

### v2

@@ 1. Introduction
Extracting insights from multiple data sources remains an arduous and time-consuming task. In fields such as biomedicine, collecting patient data requires expensive trials that typically involve only a few dozen to hundreds of patients (Li et al., 2023a). Since this usually leads to tables with many attributes but few samples, researchers need to combine different cohorts to obtain larger sample sizes. Since data is often collected using differing methods, combining them into compatible and comparable datasets often becomes a challenge (Cheng et al., 2024). Consider these examples from cancer research.
@@
To obtain a larger dataset for a study on endometrial cancer, researchers aim to combine samples from two patient cohorts collected independently in two studies (Dou et al., 2020; Dou et al., 2023). These studies produced two tables containing clinical data, which we refer to as $T_{1}$ and $T_{2}$. Each row in each table represents an individual patient sample. Tables $T_{1}$ and $T_{2}$, contain 179 and 213 columns and 153 and 190 rows, respectively. The goal is to combine the data to obtain a table with 392 rows. However, even though these datasets were produced by the same research consortium, their schemas and naming standards differ significantly. The first challenge is identifying which columns are semantically equivalent. Once all pairs of equivalent columns have been identified, their values must be standardized. Figure 1 shows an example of a pair of equivalent attributes from $T_{1}$ and $T_{2}$: Histologic_grade and Histologic_Grade_FIGO. As seen in the figure, the values of these attributes are not represented using the same terminology even though they are semantically equivalent. To produce a harmonized table, the researchers must reconcile these values into a single format before merging the rows. For instance, they may decide that the final harmonized table $T_{target}$ will contain an attribute named histologic_grade and that the format of the value used will be from Histologic_grade, and thus we would need to map the values of $T_{1}$ to their corresponding values from $T_{2}$ (e.g., "FIGO grade 1" $\rightarrow$"G1 Well differentiated", "FIGO grade 2" $\rightarrow$ "G2 Moderately differentiated", and so on). Note, however, that some attributes may be mapped differently. E.g., attributes with unique identifiers can be kept as they appear in the original table while ensuring a uniqueness constraint. ∎
@@
In a subsequent effort to foster data reuse and enable research in pan-cancer analysis (Li et al., 2023a), researchers decided to combine 10 tables containing cohorts of a larger variety of cancer types (Cao et al., 2021; Clark et al., 2019; Dou et al., 2020; Gillette et al., 2020; McDermott et al., 2020; Huang et al., 2021; Krug et al., 2020; Satpathy et al., 2021; Vasaikar et al., 2019; Wang et al., 2021a). They had to map all tables to the Genomic Data Commons (GDC) standard (Institute, 2024), a standard describing attributes commonly used in cancer research. Figure 1 shows how the variables Histologic_grade and Histologic_Grade_FIGO map to their equivalent GDC attribute tumor_grade. Once again, the acceptable values in the GDC vocabulary differ from the values used in the tables of Example 1. ∎
The Case for Agentic Data Harmonization. Data harmonization involves several data integration tasks. Practitioners often rely on spreadsheet software, bespoke scripts, and significant manual work to harmonize data (Cheng et al., 2024). These custom scripts are often not published with the data and publications, creating barriers to reproducibility and replicability of experiments with new data. (Muñoz Monjas et al., 2025).
Large language models (LLMs) open new opportunities to improve data harmonization. They can answer questions about terminology, methodologies and generate code without training data. Recently, LLMs have shown promising results in data integration tasks, including column type annotation, schema matching, and entity linkage (Narayan et al., 2022; Kayali et al., 2024; Feuer et al., 2024; Tu et al., 2023). Prompting today’s frontier LLMs with a question such as “What does FIGO grade mean?” reveals that they do learn general knowledge about many topics, including biomedical research (the focus of our examples). This suggests that this information could be leveraged in data harmonization tasks.
LLMs are becoming essential components for intelligent agents across various applications due to their capabilities in language understanding, tool utilization, and adapting to new information, resembling human intelligence and reasoning (Yao et al., 2022; Xi et al., 2023; Wang et al., 2024a). These capabilities have been surfaced by advancements in LLM prompting techniques that elicit reasoning, such as Chain-of-Thought (Wei et al., 2022), Tree-of-Thought (Yao et al., 2024), and ReAct (Yao et al., 2022). These techniques make it possible to handle complex tasks such as table understanding tasks through structured data manipulation (Wang et al., 2024b), and allow generating reasoning traces and task-specific actions that are interleaved to complete tasks. (Yao et al., 2022). Moreover, frameworks for building agentic systems, including LangChain (langchain, [n.d.]), Camel (Li et al., 2023b), Archytas (archytas, [n.d.]), and AutoGen (Wu et al., 2023), are becoming increasingly available.
@@
Second, integrating algorithms for different tasks and generating cohesive pipelines is non-trivial. Unlike in AutoML systems, data harmonization pipeline synthesis cannot be driven by search algorithms that optimize well-defined evaluation metrics (e.g., model accuracy) since the quality of data harmonization pipelines cannot be as easily measured using one metric. Instead, the synthesis may need to be guided by decisions made by the users, since accuracy may depend on external knowledge. For example, in the case of Fig. 1, deciding if the correct match for ‘NA’ is ‘Not Reported’ or ‘Unknown’ may depend on the data collection methodology. The agent should automate the laborious tasks without sacrificing accuracy and without becoming a burden: it must learn to ask questions only when necessary to avoid overwhelming the user. To add to these challenges, LLMs are brittle: (1) they often make mistakes (also known as hallucinations) that must be identified and corrected by users; and (2) they are known to be sensitive to the prompts (i.e., small prompt changes may lead to different results) (Barrie et al., 2024; Khattab et al., 2024).
@@
Contributions. This paper presents Harmonia, a prototype of an agentic data harmonization system that implements our vision. Harmonia leverages LLM capabilities to interact with users, orchestrate specialized data integration primitives, and generate custom code when existing primitives are insufficient. These primitives implement efficient and well-established data integration algorithms that can be combined to make harmonization pipelines guided based on user feedback. Since algorithms can make mistakes, we use LLMs to evaluate the outputs of these primitives. When outputs are incorrect, the agent may take additional steps to automatically correct the errors or seek assistance from the user. We also describe how Harmonia can be applied in a real-world use case, highlighting the potential and limitations of its current design and implementation. To address these limitations, we identify open problems in this field and propose future research directions, building on previous work in machine learning and data management.
In summary, we provide the following contributions:
  • We present our vision for agentic systems that help users create data harmonization pipelines by combining LLM-based reasoning, interactive user interfaces, and data integration primitives.
  • We build Harmonia, a prototype data harmonization agent that follows our vision. It integrates BDI-kit, a library of data integration algorithms, to efficiently construct harmonization pipelines. When existing functions are insufficient, Harmonia leverages the LLM to dynamically generate custom code.
  • We demonstrate a real-world use case where Harmonia is applied to map clinical data to the GDC standard (Institute, 2024), addressing common issues such as differing terminology and schema.
  • We discuss key challenges in designing data harmonization agent systems and propose a research agenda with immediate steps to address remaining open problems.
@@ 3. Agentic Data Harmonization
Figure 2. (a) An overview of the different components of an interactive, agentic data harmonization system. (b) The LLM agent loop: given a task, the LLM will repeatedly execute actions (call tools or ask questions to the user) until the task is completed. (c) A sequence diagram illustrating an example of the communication workflow between the user, agent, and primitives.
@@
We propose using LLM-based agents to facilitate the interactive construction of harmonization pipelines through natural language and visual interfaces. We aim to simplify the harmonization process and empower domain experts to harmonize their own data effectively.
@@
Our approach has three main aspects – harmonization primitives, harmonization agents, and human-agent interaction, as illustrated in Figure 2(a). The system should support two-way interaction between the users and the harmonization agents. While the former drives the system and defines the tasks to be performed, the latter aims to automate most data harmonization tasks, leaving only key decisions that need external context to users. By capturing user-agent interactions and the derived computational pipelines, we maintain the provenance of the harmonization process. This supports transparency and reproducibility, making it possible to publish the harmonized data with the pipeline used to derive it.
@@
Harmonization Primitives. The bottom of Figure 2(a) illustrates a library of components that we refer to as data integration primitives. These are key algorithms or routines that can be developed to solve well-defined data integration tasks such as schema matching, value matching, and others. Some of these components are lower-level routines that support other higher-level tasks, e.g., column-type annotation may be a component of schema-matching algorithms (Feuer et al., 2024), and entity resolution and deduplication may be key to performing data standardization (Christophides et al., 2020). While the particular set of primitives can be heterogeneous and evolve to support the system capabilities, they have in common that they need to be composable and can be invoked by both user and AI agents.
@@
Primitive composability refers to the ability to combine primitives to create a pipeline: the output of a primitive $p_{1}$ can be used as input of another primitive $p_{2}$. For example, schema matching primitives can output a list of source-target attribute pairs. These lists form the input for the value matching primitives, which then find equivalences between the values of the source attribute to the target. Finally, the output of value-matching primitives can be used to assemble a harmonization specification that describes the transformation of source tables $T_{i}$ into a target output table $T_{target}$. In short, primitive composability enables the creation of data harmonization pipelines by allowing the chaining of operations that take source tables as input and produce harmonized data as output.
@@
Data Harmonization Agents. The harmonization agent fulfills user requests by synthesizing pipelines that perform the requested harmonization task. First, it decomposes the problem in a sequence of actions. These actions can be of multiple types, including execution of existing integration primitives (tool calling) or execution of code generated on-demand. To decide on what action to take, the harmonization agent leverages an LLM, which is provided with descriptions of the task and integration primitives available for use. The LLM then returns an action (e.g., a code snippet) that can be executed in a runtime environment (e.g., a Python kernel). The action output is then fed back into the LLM, which decides if it needs to execute additional actions or if the task is completed. This loop executes until the task is deemed complete (see Figure 2(b)).
@@
Given the complexity of harmonization tasks, it is crucial to have high-level primitives available as building blocks for the pipeline. This allows encoding prior knowledge and using efficient algorithms known to be effective for a specific task. These primitives encompass algorithms for tasks such as schema matching, entity resolution, and value mapping. Of course, primitives can go beyond hard-coded functions that implement deterministic algorithms. For instance, they can be workflows that use LLMs to perform specific tasks such as in (Liu et al., 2024a) or they could generate code on demand (e.g., to extract data from or to transform attribute values (Chen et al., 2024a)).
Human-Agent Interaction. A central component of our system architecture is the user-agent interaction. We argue that systems must go beyond text-based conversational interactions: they need to support rich visual data representations and should allow interactions that help users reason about the answers produced by the agent and refine the task definition as well as the pipelines derived by the agent. In complex tasks such as harmonization, this is necessary since many decisions needed to complete the task (e.g., whether two terms represent the same concept) are difficult even for domain experts and may require external knowledge, such as the context in which data was collected.
In this paper, we focus on interactions based on natural language via a chatbot interface. However, it should be possible to implement even more intuitive and efficient graphical interfaces that combine natural language with point-and-click components. For example, as done in interactive AutoML tools (Santos et al., 2019), the system may guide the user through the harmonization process, recommend the available actions, track progress, and provide data visualizations to help the user better make sense of the data. This may help prevent common issues in natural language such as ambiguity (Esfandiarpoor and Bach, 2024).
@@ 4. System Prototype & Use Case
The primitives library. We used data harmonization primitives from bdi-kit (bdi-kit, [n.d.]), an open-source Python library that we designed with the explicit goal of composability. Currently, it includes implementations of multiple schema matching and value matching algorithms using a composable API. It includes several classic algorithms from Koutras et al. (Koutras et al., 2021) and language model-based algorithms such as the ones presented in (Liu et al., 2024b) and (Liu et al., 2024a). Most functions take as input a source parameter that represents the user’s input DataFrame and returns the output formatted as another DataFrame. The target parameter can either be a string representing a target standard schema (e.g., ‘gdc’) or a target DataFrame, this allows switching between the two tasks described in Examples 1 and  2. Figure 3 shows a subset of functions relevant to this paper.
@@
Figure 3. A subset of functions available in the bdi-kit API.
The data harmonization agent. We implemented Harmonia, a system prototype that uses the data integration primitives from Figure 3 and interacts with users via a text-based chat box. Implementing an agent entails writing carefully crafted function and task descriptions that are combined to assemble system prompts fed to the LLMs. For our prototype, we implemented tool wrappers for each of the bdi-kit functions, along with descriptions of when each should be used. We also provide general descriptions of the data harmonization steps, when the LLM should request help from the user, and output formatting instructions.
To implement the tool calling functionality and the user interface, we used the Archytas (archytas, [n.d.]) library, an open-source tool for building AI agents based on the ReAct framework (Yao et al., 2022), and Beaker (beaker, [n.d.]), a contextually-aware notebook system that supports the development of chatbot user interfaces. This allows us to implement features that facilitate interaction, such as rendering markdown-formatted outputs in the browser, displaying the reasoning steps taken by the agent, collecting user inputs, and feeding them back to the agent.
@@
[…]
Outputs:
[...data frame displayed here...]
Responding to this prompt, the agent automatically (1) generates Python code using the Pandas library that loads and selects the requested columns and (2) executes the generated code to load the CSV file into memory. This example illustrates the capability of on-demand code generation to accomplish a requested task.
@@
As shown in Figure 2(c), the system interacts with an LLM multiple times to answer this request. First, it sends the initial user’s prompt to the LLM, which recognizes that the function match_schema is adequate to solve the request, and thus returns a tool calling instruction containing the match_schema function name along with the needed parameters. Note that the LLM parses the user’s query, and recognizes that gdc must be used as the target schema and ct_learning refers to the argument method of match_schema(). After receiving the tool calling instruction as output, the system then executes the primitive function requested by the LLM, collects the list of column matches returned, and passes it back to the LLM in a follow-up prompt.
After analyzing the column match results received in the follow-up prompt, the LLM detects a mapping mistake for the column Histologic_type, which is wrongly mapped to roots. To fix the error, the LLM triggers a tool calling instruction that uses the primitive function top_matches to find the top 10 alternative candidates for Histologic_type. After receiving the results, it analyzes the results and suggests using primary_diagnosis as a correction to the user (who agrees) as seen in the following messages:
@@
As a final response to the task, the agent shows a table of column mappings as output, where the “Source Column” contains columns from the source table (dou.csv), “Target Column” contains a column match from the GDC standard, and “Corrected?” indicates whether the LLM has corrected the original bdi-kit match.
[…]
Source Column Target Column Corrected? Country country_of_birth No Histologic_Grade_FIGO tumor_grade No Histologic_type primary_diagnosis Yes, from ‘‘roots’’ FIGO_stage figo_stage No BMI bmi No Age age_at_onset No Race race No Ethnicity ethnicity No Gender gender No Tumor_Focality tumor_focality No Tumor_Size_cm tumor_depth_measurement No
Since the user is a domain expert who knows the GDC standard and knows that there may be better options for Tumor_Size_cm, it asks the system to find an alternative match. The system then updates the previous list with a more appropriate match found using the top_matches function. This interaction is shown next:
@@
As in the previous prompts, finding value mappings requires multiple LLM interactions. First, the LLM triggers one function call of match_values() for each of the source-target column pairs using the default value matching method tfidf (an efficient algorithm that chooses matches based on the tf-idf similarity over character n-grams). After receiving the value matches, the LLM evaluates if they are correct, and fixes the entries that contain errors. To do so, it calls another function to discover the list of acceptable values in a GDC attribute and picks the correct one. A snippet of the result is shown next (we omit part of the output due to lack of space):
@@
Source Value Target Value Corrected? IIIC2 Stage IIIC2 No IIIC1 Stage IIIC1 No IVB Stage IVB No IIIB Stage IIIB No IIIA Stage IIIA No II Stage II Yes, from "Stage III" [...]
@@
The final step is generating a harmonized table. The boxes below show the user prompt as well as the agent response:
@@
In this example, the agent uses materialize_mapping to generate a table and save file ‘dou_harmonized.csv’. The LLM first creates a declarative mapping specification according to the bdi-kit’s format used to describe a harmonization plan. Figure 4 shows a snippet of the generated harmonization plan, which includes a declarative specification of the mappings to be performed written in JSON. In this example, we only show dictionary-based transformations that map source values into target values. However, bdi-kit’s also supports other types of transformations, such as custom mappings that take as input a custom Python function (or lambda). This can potentially be used along with the on-demand code generation.
@@
{minted} [fontsize=]json [ […] ”source”: ”Country”, ”target”: ”country_of_birth”, ”matches”: [ [”United States”, ”United States”], [”Poland”, ”Poland”], [”Ukraine”, ”Ukraine”] ] , ”source”: ”Histologic_Grade_FIGO”, ”target”: ”tumor_grade”, ”matches”: [ [”FIGO grade 1”, ”G1”], [”FIGO grade 2”, ”G2”], [”FIGO grade 3”, ”G3”] ] , […] ]

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Novel Optimizations for Visionary AI Systems; June 22–27, 2025; Berlin, Germanybooktitle: Novel Optimizations for Visionary AI Systems (NOVAS ’25), June 22–27, 2025, Berlin, Germanydoi: 10.1145/3735079.3735324isbn: 979-8-4007-1917-2/2025/06
@@ 1. Introduction
Extracting insights from multiple data sources remains an arduous and time-consuming task. In fields such as biomedicine, collecting patient data requires expensive trials that typically involve only a few dozen to hundreds of patients (Li et al., 2023a). Since this usually leads to tables with many attributes but few samples, researchers need to combine different cohorts to obtain larger sample sizes. However, since data is often collected using differing methods, combining them into compatible and comparable datasets often becomes a challenge (Cheng et al., 2024). Consider these examples from cancer research.
@@
To obtain a larger dataset for a study on endometrial cancer, researchers aim to combine samples from two patient cohorts collected independently in two studies (Dou et al., 2020; Dou et al., 2023). These studies produced two tables containing clinical data, which we refer to as $T_{1}$ and $T_{2}$. Each row in each table represents an individual patient sample. Tables $T_{1}$ and $T_{2}$, contain 179 and 213 columns and 153 and 190 rows, respectively. The goal is to combine the data to obtain a table with 392 rows. However, even though these datasets were produced by the same research consortium, their schemas and naming standards differ significantly. The first challenge is identifying which columns are semantically equivalent. Once all pairs of equivalent columns have been identified, their values must be standardized. Figure 1 shows an example of a pair of equivalent attributes from $T_{1}$ and $T_{2}$Histologic_grade and Histologic_Grade_FIGO. The values of these attributes are represented using different terminology. To produce a harmonized table, the researchers must reconcile these values before merging the rows. For instance, they may decide that the final harmonized table $T_{target}$ will contain an attribute named histologic_grade and that the format of the value used will be from Histologic_grade, and thus we would need to map the values of $T_{1}$ to their corresponding values from $T_{2}$ (e.g., "FIGO grade 1" $\rightarrow$"G1 Well differentiated", "FIGO grade 2" $\rightarrow$ "G2 Moderately differentiated", and so on). Note that different mappings can be applied to different attribute pairs.
@@
In a subsequent effort to foster data reuse and enable research in pan-cancer analysis (Li et al., 2023a), researchers decided to combine 10 tables containing cohorts of a variety of cancer types (Cao et al., 2021; Clark et al., 2019; Dou et al., 2020; Gillette et al., 2020; McDermott et al., 2020; Huang et al., 2021; Krug et al., 2020; Satpathy et al., 2021; Vasaikar et al., 2019; Wang et al., 2021a). They mapped all tables to the Genomic Data Commons (GDC) standard (Institute, 2024), which contains attributes commonly used in cancer research. Figure 1 shows how the values for Histologic_grade and Histologic_Grade_FIGO map to the corresponding GDC attribute tumor_grade. Once again, the acceptable values in the GDC vocabulary differ from the values used in the tables of Example 1.
The Case for Agentic Data Harmonization. Data harmonization involves several data integration tasks. Practitioners use spreadsheet software, bespoke scripts, and significant manual work to harmonize data (Cheng et al., 2024). These custom scripts are often not published with the data and publications, creating barriers to the reproducibility and replicability of experiments (Muñoz Monjas et al., 2025).
Large language models (LLMs) open new opportunities to improve data harmonization. They can answer questions about terminology, methodologies and generate code without training data. Recently, LLMs have shown promising results in data integration tasks, including column type annotation, schema matching, and entity linkage (Liu et al., 2025; Kayali et al., 2024; Feuer et al., 2024; Tu et al., 2023; Narayan et al., 2022). Prompting today’s frontier LLMs with a question such as “What does FIGO grade mean?” reveals that they do have general knowledge about many topics, including biomedical research (the focus of our examples). This suggests that this information can be used in data harmonization tasks.
LLMs are becoming essential components for intelligent agents across various applications due to their capabilities in language understanding, tool utilization, and adapting to new information, resembling human intelligence and reasoning (Yao et al., 2022; Xi et al., 2023; Wang et al., 2024a). These capabilities have been surfaced by advancements in LLM prompting techniques that elicit reasoning, such as Chain-of-Thought (Wei et al., 2022) and ReAct (Yao et al., 2022). These enable agents to handle complex tasks such as table understanding tasks through structured data manipulation (Wang et al., 2024b), and generate reasoning traces and task-specific actions that are interleaved to complete tasks (Yao et al., 2022). Moreover, frameworks for building agentic systems are now widely available, such as LangChain (langchain, [n.d.]), Camel (Li et al., 2023b), and AutoGen (Wu et al., 2023).
@@
Second, integrating algorithms for different tasks and generating cohesive pipelines is non-trivial. Unlike in AutoML systems, data harmonization pipeline synthesis cannot be driven by search algorithms that optimize well-defined evaluation metrics (e.g., model accuracy) since the quality of data harmonization pipelines cannot be as easily measured using one metric. Instead, the synthesis may need to be guided by the users, since accuracy may depend on external knowledge. For example, in the case of Fig. 1, deciding if the correct match for ‘NA’ is ‘Not Reported’ or ‘Unknown’ may depend on the data collection methodology. The agent should automate the laborious tasks without sacrificing accuracy and without becoming a burden: it must learn to ask questions only when necessary to avoid overwhelming the user. To add to these challenges, LLMs are brittle: (1) they often make mistakes (also known as hallucinations) that must be identified and corrected by users; and (2) they are known to be sensitive to the prompts (i.e., small prompt changes may lead to different results) (Barrie et al., 2024; Khattab et al., 2024).
Contributions. In this paper we discuss the opportunities and challenges for the use of agents for data harmonization. We also present Harmonia, a proof-of-concept prototype of an agentic data harmonization system that implements part of our vision.
@@
Harmonia leverages LLM capabilities to interact with users, orchestrate specialized data integration primitives, and generate custom code when existing primitives are insufficient. These primitives implement efficient and well-established data integration algorithms that can be combined to make harmonization pipelines guided based on user feedback. Since algorithms can make mistakes, we use LLMs to evaluate the outputs of these primitives. When outputs are incorrect, the agent may take additional steps to automatically correct the errors or seek assistance from the user. We also describe how Harmonia can be applied in a real-world use case, highlighting the potential and limitations of its current design and implementation. To address these limitations, we identify open problems in this field and propose future research directions, building on previous work in machine learning and data management.
Our contributions can be summarized as follows: (1) We present a vision for agentic systems that help users create data harmonization pipelines by combining LLM-based reasoning, interactive user interfaces, and data integration primitives. (2) We introduce Harmonia, a prototype data harmonization agent that integrates bdi-kit (bdi-kit, [n.d.]), a library of data integration algorithms, to efficiently construct harmonization pipelines. When existing functions are insufficient, Harmonia leverages the LLM to dynamically generate custom code. (3) We demonstrate a real-world use case where Harmonia is applied to map clinical data to the GDC standard (Institute, 2024), addressing common issues such as differing terminology and schema. (4) We discuss challenges in designing data harmonization agents and propose a research agenda with steps to address open problems.
@@ 2. Preliminaries
Figure 2. (a) Components of an interactive agentic data harmonization system. Solid lines represent components implemented in Harmonia (see Section 4) while dashed lines represent components not yet implemented. (b) The LLM agent loop: given a task, the LLM will repeatedly execute actions (call tools or ask questions to the user) until the task is completed. (c) A sequence diagram illustrating an example of the communication workflow between the user, agent, and primitives.
@@ 3. Agentic Data Harmonization
We propose using LLM-based agents to facilitate the interactive construction of harmonization pipelines through natural language and visual interfaces. We aim to simplify the harmonization process and empower domain experts to harmonize their data effectively.
@@
Our approach has three main components: harmonization primitives, harmonization agents, and human-agent interaction (Figure 2(a)). The system supports a two-way interaction between users and agent. While users drive the system and define the tasks to be performed, the agent aims to automate the harmonization tasks, leaving to users only decisions that need external context. By recording user-agent interactions and the derived computational pipelines, we maintain provenance of the harmonization process, supporting transparency and reproducibility and making it possible to publish the harmonized data with the pipeline used to derive it.
@@
Harmonization Primitives. The bottom of Figure 2(a) shows a library of components that we refer to as data integration primitives. These are algorithms or routines that solve well-defined data integration tasks such as schema matching and value mapping. Some of these are lower-level routines that support other higher-level tasks, e.g., column-type annotation may be a component of schema-matching algorithms (Feuer et al., 2024), while entity resolution and deduplication may be key to performing data standardization (Christophides et al., 2020). The set of primitives can be heterogeneous and evolve to support the system capabilities. An important requirement is that they need to be composable and invoked by both user and AI agents.
@@
Composability is the ability to combine different primitives to create a data harmonization pipeline. For instance, output from schema matching primitives, which generates source-target attribute pairs, can feed into value mapping primitives that find equivalences between the values of those source and target attributes. Finally, the output of value mapping primitives can be used to create a harmonization specification that describes the transformation of source tables $T_{i}$ into a target output table $T_{target}$.
Data Harmonization Agents. An agent senses its environment and acts upon it (Russell and Norvig, 2016). In our context, we define a data harmonization agent as a program that interacts with its environment (e.g., data, primitives, users) to autonomously complete harmonization tasks. Agents can operate at various levels of autonomy (of Automotive Engineers (SAE), 2018), from offering minimal user assistance to fully independent harmonization. In Section 4, we focus on interactive agents that respond to user queries, though more autonomy is possible within our framework.
@@
To fulfill user requests, the agent must first decompose the problem into a series of actions of various types, including (1) the execution of existing integration primitives (tool calling) or (2) code generated on demand. To decompose the problem and decide on what action to take, the harmonization agent leverages an LLM, which is given descriptions of the task and integration primitives available for use. The LLM then repeatedly returns actions (e.g., a code snippet) that can be executed in a runtime environment (e.g., a Python kernel). Action outputs are fed back into the LLM, which decides if additional actions are needed. As illustrated in Figure 2(b), this loop executes until the task is deemed complete.
@@
Given the complexity of harmonization tasks, it is crucial to have high-level primitives available as building blocks for the pipeline. This allows encoding prior knowledge and using efficient algorithms known to be effective for a specific task. These primitives encompass algorithms for tasks such as schema matching, entity resolution, and value mapping. Of course, primitives can go beyond hard-coded functions that implement deterministic algorithms. For instance, they can be workflows that use LLMs to perform specific tasks such as in (Liu et al., 2025) or they could generate code on demand (e.g., to extract data from or to transform attribute values (Chen et al., 2024a)).
Human-Agent Interaction. A key aspect of our architecture is the interaction between users and agents. In Section 4, we show that our current prototype is proficient in text-based conversational interactions: it can parse users’ input and act on them. However, to make systems more effective, we argue that future systems need to support graphical user interfaces that help users (1) to reason about the input data and the agent’s output, and (2) to refine the task definition and pipelines derived by the agent. For example, as done in interactive AutoML tools (Santos et al., 2019), the system could guide the user through the harmonization process, recommend the available actions, track progress, and provide data visualizations to help the user better make sense of the data. This may help prevent common issues in natural language such as ambiguity (Esfandiarpoor and Bach, 2024; Yu and Silva, 2019).
In complex tasks like harmonization, user interaction is crucial because many decisions—such as whether two terms represent the same concept—can be difficult even for domain experts. These judgments often depend on context not captured in the data, such as how it was collected or the purpose of the harmonization. Furthermore, as in any automated data integration system, LLM-based agents can make mistakes. When combined with well-designed user interfaces, users are better equipped to guide the agent, providing guardrails and safety mechanisms that neither users nor interfaces could achieve as effectively on their own.
@@ 4. System Prototype & Use Case
We implemented a proof-of-concept of our vision (Figure 2(a)). In what follows, we illustrate how it works with a concrete example.
Data Integration Primitives. We used data harmonization primitives from bdi-kit (bdi-kit, [n.d.]), an open-source Python library that we designed with the explicit goal of composability. Currently, it includes implementations of multiple schema matching and value mapping algorithms using a composable API. It also includes several classic algorithms from Koutras et al. (Koutras et al., 2021) and language model-based algorithms (Liu et al., 2024, 2025). Most functions take as input a source parameter that represents the user’s input DataFrame and returns the output formatted as another DataFrame. The target parameter can either be a string representing a target standard schema (e.g., ‘gdc’) or a target DataFrame, this allows switching between the two tasks described in Examples 1 and  2. Figure 3 shows a sample of functions integrated in Harmonia. The source code is on GitHub (Harmonia, [n.d.]).
@@
Figure 3. Some bdi-kit functions integrated in Harmonia.
LLM Agent & User Interface. Implementing an agent entails writing carefully crafted function and task descriptions that are combined to assemble system prompts fed to the LLMs. For our prototype, we implemented tool wrappers for each of the bdi-kit functions, along with descriptions of when each should be used. We also provide descriptions of the data harmonization steps, when the LLM should request help from the user, and output formatting instructions. These descriptions are available in the code (Harmonia, [n.d.]).
To implement tool calling, we used the Archytas (archytas, [n.d.]), an open-source library for building AI agents based on the ReAct framework (Yao et al., 2022), and used the GPT-4o model (i.e., no fine-tuning is involved). Instead of using a fixed pipeline that invokes the LLM, we implement a dynamic agentic loop, as described in Figure 2(b). The LLM has the flexibility to break down the task into multiple tool call instructions needed to complete the task. The user interface is implemented using Beaker (beaker, [n.d.]) and is currently limited to a chat-based interface and workflow that resembles a Jupyter notebook. The system renders markdown-formatted outputs in the browser, displays the reasoning steps taken by the agent, and collects user inputs using text input boxes and feeds them back to the agent. A video demonstrating these interactions is available online (Harmonia Demonstration., [n.d.]).
@@
As a response, the agent shows a success message and a table preview. To produce this result, the agent automatically (1) generates Python code using the Pandas library that loads and selects the requested columns and (2) executes the generated code to load the CSV file into memory. This example illustrates the capability of on-demand code generation to accomplish a requested task.
@@
As shown in Figure 2(c), the system interacts with an LLM multiple times to answer this request. First, it sends the initial user’s prompt to the LLM, which recognizes that the function match_schema is adequate to solve the request, and thus returns a tool calling instruction containing the match_schema function name along with the needed parameters. After receiving the instruction, the system executes the function, collects matches, and returns them to the LLM. Note that the LLM interprets the user’s query and correctly identifies the function arguments: it uses gdc as the target schema and a default value for the method parameter in match_schema(). Alternatively, the agent also supports requesting a specific method in the prompt, offering experts greater control and customization.s
After analyzing the column match results received in the follow-up prompt, the LLM detects a mapping mistake for the column Histologic_type, which is incorrectly mapped to roots. To fix the error, the LLM triggers a tool calling instruction that uses the primitive function top_matches to find the top-10 alternative candidates for Histologic_type. After analyzing the results, it suggests using primary_diagnosis as a correction to the user.
@@
This exemplifies an important benefit of the agent in augmenting and improving the effectiveness of the underlying libraries: using its knowledge, it is able to detect the error; then, it leverages the library to identify alternative choices. This also shows an example of complex task planning enabled by the agentic loop (Figure 2(b)).
The agent then displays a table of column mappings, where the “Source Column” contains columns from the source table (dou.csv), “Target Column” contains a column match from the GDC standard, and “Corrected?” indicates whether the LLM has corrected the original bdi-kit match.
@@
Source Column Target Column Corrected? [...] [...] [...] Country country_of_birth No Histologic_Grade_FIGO tumor_grade No Histologic_type primary_diagnosis Yes, from ‘‘roots’’ Tumor_Size_cm tumor_depth_measurement No
Since the user is a domain expert familiar with the GDC standard, they ask for better matching options for Tumor_Size_cm. The system updates the previous list with a more appropriate match found using the top_matches function:
@@
As in the previous prompts, finding value mappings requires multiple LLM interactions. First, the LLM call the match_values() function for each of the source-target column pairs using the default method tfidf (an efficient algorithm that chooses matches based on the tf-idf similarity over n-grams). After receiving the value mappings, the LLM verifies if they are correct and fixes the erroneous entries. To do so, it calls another function to discover the list of acceptable values in a GDC attribute and picks the correct one. A snippet of the result is shown next (we omit part of the output due to lack of space):
@@
Source Value Target Value Corrected? IIIB Stage IIIB No IIIA Stage IIIA No II Stage II Yes, from "Stage III"
@@
The final step is to generate a harmonized table:
@@
Here, the agent uses materialize_mapping to generate a table and save file ‘dou_harmonized.csv’. The LLM first creates a declarative mapping specification according to the bdi-kit’s format used to describe a harmonization plan. Figure 4 shows a snippet of the generated harmonization plan, which includes a declarative specification of the mappings in JSON. In this example, we only show dictionary-based transformations that map source values into target values. However, bdi-kit also supports other types of transformations, such as custom mappings that take as input a custom Python function (or lambda). This can potentially be used along with the on-demand code generation.
@@
[ […] { source”: Histologic_Grade_FIGO”, target”: tumor_grade”, matches”: [ [”FIGO grade 1”, G1”], [”FIGO grade 2”, G2”], [”FIGO grade 3”, G3”] ] }, […] ]
@@
Evaluation. We conducted a preliminary evaluation and compared Harmonia against baseline methods from bdi-kit executed without agent support. The results, summarized in Table 1, show that Harmonia achieved the best performance across both tasks. For schema matching, Harmonia successfully associated the column ‘Histologic_type’ with ‘primary_diagnosis’, which contains 2,625 unique values in the GDC. For value mapping, Harmonia was able to correct values such as ‘FIGO grade 1’ to ‘G1’.
Table 1. Performance comparison for schema matching and value mapping tasks.
Task Method Accuracy Precision Recall F1
Schema Baseline 0.88 0.78 0.78 0.78
Matching Harmonia 1.00 1.00 1.00 1.00
Value Baseline 0.58 0.58 0.59 0.57
Mapping Harmonia 0.68 0.69 0.69 0.68